OpenStudy (lostinmath101):

Find the inverse of the given function and use composition to justify the results: h(x)=(x-4)^3

OpenStudy (anonymous):

to "justify your answer" compose \[f\circ f^{-1}(x)\] and arrive at \(f\circ f^{-1}(x)=x\) by algebra

OpenStudy (lostinmath101):

Were did you get the 3 from?

OpenStudy (anonymous):

the three came from me not paying sufficient attention it is \[f^{-1}(x)=\sqrt[3]{x}+4\] sorry about that

OpenStudy (anonymous):

\[y=(x-4)^4\] \[x=(y-4)^3\] \[\sqrt[3]{x}=y-4\] \[\sqrt[3]{x}+4=y\] \[f^{-1}(x)=\sqrt[3]{x}+4\] looks a lot better

OpenStudy (lostinmath101):

Thanks